OpenStudy (happykiddo):

The magnitude of displacement a and b are 3 m and 4 m respectively. Considering various orientation of a and b, what is i) maximum magnitude for c and ii) the minimum possible magnitude?

OpenStudy (happykiddo):

This is vectors, and I just cant understand how c-min=a-b=3-4=1 instead of -1 ?????